(4) (15 marks) Repeat the Question 2 for the following matrices -3 4 0] 0 0 A -2 30 B 0 -1 0 -8 8 1 0 0 1 ū= 10 = > 3(I) (2 mark) Find the characteristic polynomial of matrix A. (II) (1 mark) Find eigenvalues of the matrix A. III) (2 mark) Fi
Answer #1

-3 4 0 A= -2 o 3 8. -8 (I we have I = 1 об 0 to P(x) = det CA-XI) where P() is the characteristic polynanial of il matrix A (& x²(x-1)-1(1-1) 20 (+²1) (x-1)=0 -- | |-رارا = ۸ X eigen values of A. (III) Bacis for the eigenspaces of matrix A. we have =If we take X₂=t, z = 5 then 2,=t, d2 = t , & = 5 Therefore, + x= Itt lo S. t Lot S 19:30 eigen rectors are Hence, the 1 ero 6(A + I)x=0 SPA-AND Now o 0 -Ź HE O &z=t then Az If we take 2,2 t , Therefore, zt 4 Azt 2 x = th to t - 4 Hence elgen vector iNon G.M for x=1 is [A-I] .4 4 o -2 2 - 8 Here the rank of this is 2 at so G.M = n- Rank 2 3 - 2 - 1 For x=1 - G.M= 1 For x=-1-3 D A = 3 -8 8 1 Matrix A is diagonalizable because its all eigen vectors are different and exist eigon vector are :- 0 and